# IAP: Improving Continual Learning of Vision-Language Models via Instance-Aware Prompting

### v1

@@ 4.2 Experimental Results
Figure 5: The process of our IA-CDDP module for test samples across tasks 1, 4, and 9. As illustrated in the first row of the figure, the first stage of IA-CDDP constructs distributions at the task level and performs inference based on the similarity between samples and each distribution. Test samples with either excessively low or high confidence scores are directly classified, while samples with intermediate confidence scores proceed to the second stage. At this stage, we construct distributions for all classes within the task and calculates similarities with the 5 nearest class distributions, highlighted by the red dashed circle in the second row. The mean of these distribution similarities serves as the confidence score for the test samples, which subsequently guides the following prompting process.
Trainable parameters and memory buffer. We present a comparison of the trainable parameters and rehearsal memory buffer requirements for our method and four other SOTA approaches in Table 2. Leveraging PEFT strategies, our method eliminates the need for external storage of representative samples or features. Furthermore, our approach utilizes only $1/25$ of the trainable parameters compared to the MoE-Adapter [47] method. This substantial reduction is achieved because our method does not require the training of numerous experts; instead, for each task, only the corresponding prompt pool and prompt gate module are trained.
@@
Effect of Instance-Aware Gated Prompting strategy. We conducted a statistical analysis of mean prompting layers employed by our IA-GP strategy across different distributions, as illustrated in Figure 4. The red line represents the prompt addition strategy employed by DIKI [36], which consistently prompting the first eight transformer layers. In contrast, our IAP method, benefiting from the IA-GP strategy, demonstrates greater flexibility by dynamically determining whether to apply prompts to different transformer layers at the instance level. Experimental results indicate that for tasks with more complex distributions, such as SUN397 (397 classes) and Food101 (101 classes), the IA-GP strategy tends to incorporate more prompts to effectively capture intricate distributional information. Conversely, for relatively simpler datasets like DTD (47 classes) and EuroSAT (10 classes), the strategy selectively employs fewer prompts, thereby achieving superior performance while reducing inference time. We also provide the specific prompting layers for each task, presented as a heatmap. For details, please refer to the supplementary materials.
@@
Effect of Instance-Aware Class-Distribution-Driven Prompting strategy. In Figure 5, we visualize the distributions of all 11 tasks and the class distributions of three specific tasks: Aircraft (100 classes), OxfordPet (47 classes), and StanfordCars (196 classes). For each task, we use a corresponding instance as an example, denoted by a star. The reference distance, displayed in the lower left corner of each graph, serves as an indicator of confidence score, with values closer to 1 signifying higher confident. The top row of the figure illustrates the distributions of the 11 tasks. The DIKI [36] method leverages the similarity between an instance and the tasks distribution to derive confidence scores, which in turn control the weight of subsequent prompting operation. In contrast, our IA-CDDP method incorporates a second stage that computes the average of the five nearest class distributions (highlighted by red dotted circles) at the instance level. The reference distances demonstrate that our IA-CDDP method assigns more appropriate instance-aware confidence scores, thereby enhancing the efficacy of subsequent prompts.
@@
Quantitative Analysis Our method is implemented on a classic prompt-based MCIL method DIKI [36]  [^1] https://github.com/lloongx/DIKI. We perform a quantitative analysis of our proposed IA-GP and IA-CDDP modules and the results are shown in Table 3. For the baseline method, we observed that the reproduced results (DIKI) are lower than the papers’ results (DIKI*). Compared to baseline DIKI, both IA-GP and IA-CDDP can achieve a performance improvement. The details are included in the supplemental material.
5 Conclusion
@@
In this paper, we propose the Instance-Aware Prompting method to tackle the challenges of backward forgetting and forward forgetting faced by pre-trained VLMs during incremental learning of new tasks. The IAP method comprises two meticulous strategies designed at the instance level. The Instance-Aware Gated Prompting strategy dynamically controls whether to prompt at the instance level by incorporating a prompt gate module for each transformer layer. The Instance-Aware Class-Distribution-Driven Prompting strategy assigns more precise weights to each instance through a two-stage determination process. By integrating these two strategies, our method simultaneously resolves the questions of whether to apply prompt and how much prompt to apply, enabling pre-trained VLMs to adapt more flexibly when incrementally learning new tasks. Extensive experimental results demonstrate that the IAP method represents the current SOTA approach in the MCIL benchmark.

### v2

@@ IV-B5 Instance-Aware Gated Baselines
Figure 5: The classification process of our IA-CDDP module for test samples across tasks 1, 4, and 9. As illustrated in the first row of the figure, the first stage of IA-CDDP constructs distributions at the task level and performs inference based on the similarity between samples and each distribution. Test samples with either excessively low or high confidence scores are directly classified, while samples with intermediate confidence scores proceed to the second stage of IA-CDDP. At this stage, we construct distributions for all classes within the task and calculates similarities with the 5 nearest class distributions, highlighted by the red dashed circle in the second row. The mean of these distribution similarities serves as the confidence score for the test samples, which subsequently guides the following prompting process.
IV-B6 Effect of Instance-Aware Gated Prompting Strategy
@@
We conduct a statistical analysis of the mean prompting layers employed by our IA-GP strategy across different distributions, as illustrated in Figure 4. The red line represents the prompt injection strategy employed by DIKI [9], which consistently prompting the first 8 layers. In contrast, our method, benefiting from the IA-GP strategy, demonstrates greater flexibility by adaptively determining whether to apply prompts to different transformer layers at the instance level. Experimental results indicate that for tasks with more complex distributions, such as SUN397 (397 classes) and Food101 (101 classes), the IA-GP strategy tends to incorporate more prompts to effectively capture intricate distributional information. Conversely, for relatively simpler datasets like DTD (47 classes) and EuroSAT (10 classes), the strategy employs fewer prompts, thereby achieving superior performance while avoiding overfitting.
IV-B7 Effect of Instance-Aware Class-Distribution-Driven Prompting Strategy
@@
In Figure 5, we visualize the distributions of all 11 tasks and the class distributions of three specific tasks: Aircraft (100 classes), DTD (47 classes), and OxfordPet (37 classes). For each task, we use a corresponding instance as an example, denoted by a star. The reference distance, displayed in the lower left corner of each graph, serves as the confidence score, with values closer to 1 signifying higher confident. The top row of the figure illustrates the distributions of the 11 tasks. The DIKI [9] method leverages the similarity between an instance and the tasks distribution to derive confidence scores. In contrast, our IA-CDDP strategy incorporates a second stage that computes the average of the 5 nearest class distributions (highlighted by red dotted circles) at the instance level. The reference distances demonstrate that our IA-CDDP strategy assigns more appropriate instance-aware confidence scores, thereby enhancing the efficacy of subsequent prompts. We also conduct ablation study on the bounds of IA-CDDP strategy, the details are provided in the supplemental materials.
IV-B8 Quantitative Analysis
@@
Our method is implemented on a classic prompt-based MTIL method DIKI [9]. We perform a quantitative analysis of our proposed IA-GP and IA-CDDP strategies and the results are shown in Table V. For the baseline method, we observed that the reproduced results (DIKI) are lower than the papers’ results (DIKI*). Compared to baseline DIKI, both IA-GP and IA-CDDP can achieve a performance improvement.
V Conclusions
@@
In this paper, we propose the Instance-Aware Prompting method to solve the challenges of backward forgetting and forward forgetting faced by pre-trained VLMs during incremental learning of new tasks. Our method comprises two meticulous strategies designed at the instance level. The Instance-Aware Gated Prompting strategy dynamically controls whether to prompt at the instance level by incorporating a prompt gate module for each transformer layer. The Instance-Aware Class-Distribution-Driven Prompting strategy assigns more precise weights to each instance through a two-stage determination process. By integrating these two strategies, our method simultaneously resolves the questions of whether to apply prompt and how much prompt to apply, enabling pre-trained VLMs to adapt more flexibly when incrementally learning new tasks. Extensive experimental results demonstrate that the IAP method represents the current SOTA approach in the MTIL benchmark.